OpenStudy (anonymous):

What is the local minimum value of the function? g(x)=x⁴-5x²+4

OpenStudy (anonymous):

you need to differentiate the fuction and set it equal to zero. \[g \prime(x)=4x ^{3}-10x\]\[0=4x ^{3}-10x\]\[x=\sqrt{5/2}\] so \[g(\sqrt{5/2})\]should give you the answer

OpenStudy (anonymous):

-9/4 is the answer in this case

OpenStudy (cwrw238):

differentiate and equate to 0: 4x^3 - 10x = 0 2x(2x^2 - 5)=0 x = 0 or +- sqrt (5/2) so there re turning points at the x values above find second derivative = 12x^2 - 10 now plug in the above values of x into second derivative - the positive value will give you the local minimum

OpenStudy (anonymous):

yup zeros are sqrt(5/2), -sqrt(5/2) and 0

OpenStudy (anonymous):

and then if I round to the nearest hundredth

OpenStudy (anonymous):

-2.25

OpenStudy (anonymous):

first you have to factorise it like this lets say x^2=a \[a^2-5a+4=0\] \[(a-4)(a-1)=0\] now you can write x^2 to places where you see a \[(x-2)(x+2)(x-1)(x+1)=0\] this is our equation than try to make this minimum by giving values

OpenStudy (cwrw238):

local minimums are at x= sqrt(5/2), x= - sqrt(5/2)

OpenStudy (anonymous):

yeah but it is an even function so both of them will give the same value thats why i didnt specify :)

OpenStudy (cwrw238):

yea - no probs